# Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge

### v1

@@ Abstract
$H^{\prime}=H+\sum_{\ell}h_{x}X_{\ell},$ (8)
with an additional transverse field which breaks both U(1) symmetry and integrability. When breaking U(1) symmetry, we use the parameters $h_{x}=0.75$ and $J_{2}=0$.
We now analyze the EE distribution of mid-spectrum eigenstates at half-filling as a function of $L$ and $f$. In Fig. 2(a), we show the EE distribution of mid-spectrum eigenstates as a function of $L$ for fixed $f=1/2$, both for the Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) with U(1) symmetry (circles) and for the Hamiltonian (8) without U(1) symmetry (squares). Each data point represents the average EE of eigenstates relative to the maximum entanglement entropy $L_{A}\log(2)$, and the bars indicate their standard deviation. The shaded blue area indicates the regions limited by $\mu_{M}\pm\sigma_{M}$ for the RMT ensemble $|\Psi\rangle\in{\cal H}(M)$ using real-valued vectors, and the shaded red area indicates the region limited by $\mu_{M,N}\pm\sigma_{M,N}$ for the ensemble $|\Psi\rangle\in{\cal H}(M,N)$ also using real-valued vectors (the boundaries are interpolated smoothly between the accessible $L$ values). We obtain the eigenstate distributions by choosing an energy window, $W=20,50,100,600$ of eigenstates for $L=10,12,14,16$ around the peak density-of-states. The entanglement bipartition is placed in the center of the chain to prevent edge effects from the open boundary conditions. For the first and second moments of the EE distribution, we find excellent quantitative agreement between the eigenstate distribution and the corresponding constrained RMT ensemble. This applies both for the Hamiltonian with a U(1) scalar charge and the Hamiltonian without additional structure.
Figure 3: Finite-size scaling of the average EE of mid-spectrum eigenstates relative to the RMT prediction $\mu_{M,N}$ plotted as a function of the integrability breaking parameter $J_{2}$. The inset shows the standard deviation of EE of mid-spectrum eigenstates relative to the RMT prediction $\sigma_{M,N}$. Excellent agreement with the constrained RMT ensemble (at the level of the first two moments) is found only in a small pocket of model parameters centered around $J_{2}\approx 0.6$. Parameters used: $J=1$, $\Delta=J/2$, $h_{x}=0$, $f=1/2$.
@@
Whereas previous works have made important steps in highlighting deviations from maximal entropy of eigenstates in Hamiltonian systems with a U(1) charge[19, 43], with deviations qualitatively captured by the ensemble $|\Psi\rangle\in{\cal H}(M)$, clear differences between the eigenstate and the U(1)-constrained RMT ensembles are visible when both ensembles are compared on the microscopic scale given by the entropy fluctuations $\sigma_{M}$. In particular, we find that the typical EE of eigenstates in Hamiltonian systems with a U(1) charge is several standard deviations away from the U(1)-constrained RMT ensemble, as shown in Figs.1 and 2. Overall, the presence of two local charges—energy and magnetization—effectively constrains the available phase space, leading to quantitatively distinct patterns of entanglement at the level of first and second moments.
Maximally chaotic systems.—In Figs. 1 and 2, we set the next nearest-neighbor term $J_{2}=0.6$, which we found to be the value where eigenstates are most random. In Ref. [35], we conjectured the existence of ‘maximally chaotic’ Hamiltonians at fined tuned regions of parameter space where the distance to the appropriately constrained RMT ensembles is minimal. The model used in Ref. [35] was the Mixed Field Ising Model, with the higher moments of the EE minimizing and agreeing with RMT ensembles in small pockets of parameter space. We reach the same conclusion in a completely different model, as shown in Fig. 3. For instance, by tuning $J_{2}$ away from $J_{2}=0.6$ [vertical dashed line], we find systematic deviations from the constrained RMT ensembles that persist in the thermodynamic limit. In particular, eigenstates and constrained RMT states exhibit the same patterns of entanglement at the level of first and second moments when $J_{2}\approx 0.6$. As we depart from $J_{2}\approx 0.6$, there is a decrease in the average entanglement and an increase in their fluctuations that persist in the thermodynamic limit.
@@
Discussion.—Our work shows that constrained RMT ensembles can describe increasingly finer features of eigenstates in physical systems, with key system properties captured by appropriate constraints. This sharpens our understanding of quantum chaos and its connection with RMT. This also shows the need to characterize new classes of constrained RMT ensembles to understand the universal properties of eigenstates in physical systems with additional symmetries. An interesting direction for future research is describing the fine features of EE in the presence of non-abelian symmetries [44, 45]. On a different front, our work strengthens our claim in Ref. [35], suggesting a notion of ‘maximal chaos’ in local physical systems in fined-tuned pockets of parameter space. In particular, ensembles of eigenstates are indistinguishable from appropriately-constrained RMT ensembles at the level of subsystem entropies only in small pockets of parameter space. An interesting question is, what conditions are required to maximize chaos in local systems, and what are the dynamical signatures of these ‘maximally chaotic’ Hamiltonians?
@@
Acknowledgements.—We thank Vedika Khemani and Cheryne Jonay for previous collaborations and feedback on the present work, as well as Nick Hunter-Jones and Sam Garratt for helpful discussions. JFRN acknowledges the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and where this work was initiated. The numerical simulations in this work were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.
@@
${}^{1}$epartment of Physics & Astronomy, Texas A&M University, College Station, TX 77843
This Supplementary Material discusses the analytical and numerical details of the results presented in the main text. Section S1 discusses the entanglement entropy (EE) distributions generated by constrained and unconstrained random matrix theory (RMT) ensembles. Section II focuses on the statistical analysis of the numerical data.
@@
In this section, we consider the statistical properties of the EE for various RMT ensembles, both constrained and unconstrained. We start by discussing the case of pure random states without constraints, and then discuss the cases with U(1) and U(1)$\times$×U(1) constraints. In all cases, we discuss the exact analytical results as well as the asymptotic behavior, which were discussed in the main text. We finally discuss differences between ensembles drawn from complex valued distributions (GUE) and real valued distributions (GOE).
@@
$\rho_{A}$ $=\sum_{\alpha,\alpha^{\prime}=1}^{d_{A}}\left[\sum_{\beta=1}^{d_{B}}\psi_{\alpha,\beta}\psi^{*}_{\alpha^{\prime},\beta}\right]\ket{\alpha}\bra{\alpha^{\prime}}.$ | . (S2)
@@
We now discuss the distribution of EE of pure random states with a U(1)$\times$ U(1) constraint, defined by a Hilbert space, $\mathcal{H}(M,N)$ with a fixed number of particles $M$ and $N$. Let us consider a system of $L$ sites with a local Hilbert space dimension $d=4$, with the local basis states spanned by either an empty site, a site occupied by particle-$1(2)$ or occupied by both particles [see Fig. (1)(b) in the main text]. The Hilbert space dimension of a symmetry sector with $(M,N)$ particles is
@@
with $\sum_{M,N=0}^{L}\binom{L}{M}\binom{L}{N}=4^{L}$ by Vandermonde’s identity. We next bipartition the system into $A$ and $B$ subregions of size $L_{A}$ and $L_{B}=L-L_{A}$. In this case, the Hilbert space of the $(M,N)$ symmetry sector decomposed as
@@
where the second term is the Shannon entropy which captures classical correlations between particles of the subsystems, while the first term captures quantum correlations. Importantly, since the two $U(1)$ charges commute, we assume that $p_{M_{A},N_{A}}=p_{M_{A}}p_{N_{A}}$ are independent probabilities for each flavor. In addition, each block $\rho_{A|M_{A},N_{A}}$ with $(M_{A},N_{A})$ particles is assumed to be a random matrix. In this case, $S(\rho_{A|M_{A},N_{A}})$ is the Page entropy (S3) for pure random strained constrained to $(M_{A},N_{A})$ particles in subsystem $A$. Thus, the average EE $\mu_{M,N}=\langle S_{A}\rangle_{M,N}$ for pure random states in the $(M,N)$ symmetry sector is the Page entropy for each block $\rho_{A|M_{A},N_{A}}$ averaged with the distribution $p_{M_{A},N_{A}}$,
@@
Unlike the variance $\sigma_{M}^{2}$ in Eq.(S19), however, here we note that $\sigma_{M,N}^{2}$ is a factor ${L}$ larger than $d^{L}$ due to the midspectrum symmetry sector $d_{M,N}$ being a factor $1/L$ smaller than the dimension of the full Hilbert space. As such, we find that that $\sigma_{M,N}/\sigma_{M}\sim L^{1/4}$, see Fig.S1(d).
@@
In Fig. (S1), we numerically compare the first two moments of the entanglement distribution for different system sizes using the random states in Eq. (S1.3) where the $\psi_{\alpha,\beta}^{(M_{A},N_{A})}$ are drawn either from a GUE or GOE ensemble. While exact expressions for the first moments differ between GUE and GOE, they are asymptotically the same in the thermodynamic limit [39, 46, 42]. Fig. S1(a) confirms that there is little difference between the average entanglement values. Figure S1(b) also shows that the standard deviation of $\sigma_{M,N}$(GUE) is a factor $\sqrt{2}$ larger than $\sigma_{M,N}$(GOE) (shown by the blue dashed line). The additional prefactor is due to averaging over both the real and complex parts of the GUE ensemble.
@@
For the Hamiltonian in Eq.(Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) of the main text, when computing the first two moments of the EE distribution for mid-spectrum eigenstates, it is necessary to take a finite energy window $\Delta E$ in which to take samples of $S_{A}$, see Fig. S2(a). If the energy window is too small, then a statistically small number of states will be available for sampling, thus leading to large error bars. On the other hand, if the window is too large, then finite-energy eigenstates with low entanglement will skew the distribution and increase its variance, see Figs.S2(c). Because of typicality, we argue only a few eigenstates are necessary to quantify the mean and standard deviation. In Fig. S2 (a) we plot the eigenstate entanglement of the full spectrum for different system sizes, highlighting the set of eigenstates used to construct the microcanonical average [indicated with vertical lines in (b) and (c)].

### v2

@@ Abstract
$H^{\prime}=H+h_{x}\sum_{\ell}X_{\ell},$ (9)
with an additional transverse magnetic field which breaks both U(1) symmetry and integrability. When breaking U(1) symmetry, we use $h_{x}=0.75$ and $J_{2}=0$.
We now analyze the EE distribution of mid-spectrum eigenstates at half-filling as a function of $L$ and $f$. In Fig. 2(a), we show the EE distribution of mid-spectrum eigenstates as a function of $L$ for fixed $f=1/2$, both for the Hamiltonian (Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) with U(1) symmetry (circles) and for the Hamiltonian (9) without U(1) symmetry (squares). Each data point represents the average EE of eigenstates relative to the maximum entanglement entropy $L_{A}\log(2)$, and the bars indicate their standard deviation. The shaded blue area indicates the regions limited by $\mu_{M}\pm\sigma_{M}$ for the ensemble $|\Psi\rangle\in{\cal H}(M)$ using real-valued vectors, and the shaded red area indicates the region limited by $\mu_{M,N}\pm\sigma_{M,N}$ for the ensemble $|\Psi\rangle\in{\cal H}(M,N)$ also using real-valued vectors (the boundaries are interpolated smoothly between the accessible $L$ values). We obtain the eigenstate distributions by choosing an energy window, $W=20,50,100,600$ of eigenstates for $L=10,12,14,16$ around the peak density-of-states. The entanglement bipartition is placed in the center of the chain to prevent edge effects from the open boundary conditions. For the first and second moments of the EE distribution, we find excellent quantitative agreement between the eigenstate distribution and the corresponding constrained ensemble. This applies both for the Hamiltonian with a U(1) scalar charge and the Hamiltonian without additional structure.
Figure 3: Finite-size scaling of the average EE of mid-spectrum eigenstates relative to the analytical prediction $\mu_{M,N}$ in Eq.(7) plotted as a function of the integrability breaking parameter $J_{2}$. The inset shows the standard deviation of EE of mid-spectrum eigenstates relative to the analytical prediction $\sigma_{M,N}$. Excellent agreement with the constrained ensemble (at the level of the first two moments) is found in a small pocket of model parameters centered around $J_{2}\approx 0.6$. Parameters used: $J=1$, $\Delta=J/2$, $h_{x}=0$, $f=1/2$.
@@
As a final remark, we compare our results with previous works[17, 19, 45] that have also looked at the behavior of eigenstate EE in Hamiltonian systems with an additional U(1) charge, such as that described in Eq.(Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge). These previous works found excellent agreement between the EE distribution of eigenstates and that of random state ensembles with a single U(1) charge away from half-filling. However, they observed deviations at half-filling that could not be explained by the same ensemble. These deviations align with our findings in Fig. 1, which show large discrepancies when comparing the EE distribution of eigenstates relative to that of U(1)-constrained random states. We argue that the excellent agreement away from half-filling arises because, in this regime, a sizable $O(\sqrt{L})$ correction dominates over the $O(1)$ term [18]. In contrast, at half-filling, the $O(\sqrt{L})$ term vanishes, making the $O(1)$ term the dominant contribution. In this case, it becomes important to appropriately constrain the states with two U(1) scalar charges—one for energy and the other for magnetic charge.
Maximally chaotic systems.—In Figs. 1 and 2, we set the next-nearest-neighbor coupling to $J_{2}=0.6$, as we found this value to produce the most random eigenstates. We now extend our analysis to a broader range of $J_{2}$ values, as shown in Fig. 3. Consistent with the results in Ref.[35] for the Mixed Field Ising Model, we observe remarkably-good agreement between the eigenstate ensemble and the (appropriately) constrained ensemble of random states specifically in small regions of parameter space near $J_{2}\approx 0.6$, even for relatively small system sizes ($L=12$). Outside these regions, we find finite-size deviations between the ensembles which vanish in the thermodynamic limit. These findings support the conjecture in Ref.[35] that maximally chaotic Hamiltonians exist in small pockets of parameter space, where even for modest system sizes, the distance to the appropriately constrained ensembles is minimized.
@@
Discussion.—Our work shows that constrained ensembles of pure random states can accurately describe increasingly finer features in the eigenstate statistics of physical systems, with the constraints capturing the essential features of the Hamiltonian, namely spatial locality and conservation laws. This also highlights the need to characterize new classes of random state ensembles to understand the universal properties of eigenstates in other physical systems, such as those with non-abelian symmetries [46, 47, 48] or with integrable limits[19, 49, 50, 51]. On a different front, our work strengthens the claim in Ref. [35], suggesting a notion of ‘maximal chaos’ emerging in local Hamiltonians. Interesting questions to address are, what conditions are required to maximize chaos in local systems, and what are the dynamical signatures of these ‘maximally chaotic’ Hamiltonians?
@@
Acknowledgements.—We thank Vedika Khemani and Cheryne Jonay for previous collaborations and feedback on the present work, as well as Thomas Barthel, Nick Hunter-Jones, Sam Garratt, and Marcos Rigol for insightful comments. JFRN acknowledges the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and a grant from the Alfred P. Sloan Foundation (G-2024-22395). The numerical simulations in this work were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.
@@
1Department of Physics & Astronomy, Texas A&M University, College Station, TX 77843
This Supplementary Material discusses the analytical and numerical details of the results presented in the main text. Section S1 discusses the entanglement entropy (EE) distributions generated by constrained and unconstrained random state ensembles. Section II focuses on the statistical analysis of the numerical data.
@@
In this section, we consider the statistical properties of the EE for different ensembles of random states, both constrained and unconstrained. We start by discussing the case of pure random states without constraints, and then discuss the cases with U(1) and U(1)$\times$×U(1) constraints. In all cases, we discuss the exact analytical results as well as the asymptotic behavior, which were discussed in the main text. We finally discuss differences between ensembles drawn from complex valued distributions (GUE) and real valued distributions (GOE).
@@
$\rho_{A}$ $|.\displaystyle=\sum_{\alpha,\alpha^{\prime}=1}^{d_{A}}\left[\sum_{\beta=1}^{d_{B}}\psi_{\alpha,\beta}\psi^{*}_{\alpha^{\prime},\beta}\right]\ket{\alpha}\bra{\alpha^{\prime}}.$ | . (S2)
@@
We now discuss the distribution of EE of pure random states with a U(1)$\times$ U(1) constraint, defined by a Hilbert space, $\mathcal{H}(M,N)$ with a fixed number of particles $M$ and $N$. Let us consider a system of $L$ sites with a local Hilbert space dimension $d=4$, with the local basis states spanned by either an empty site, a site occupied by particle-1(2)1(2)1 ( 2 ) or occupied by both particles [see Fig. (1)(b) in the main text]. The Hilbert space dimension of a symmetry sector with $(M,N)$ particles is
@@
[…]
@@
[…]
@@
[…]
@@
In Fig. (S1), we numerically compare the first two moments of the entanglement distribution for different system sizes using the random states in Eq. (S1.3) where the $\psi_{\alpha,\beta}^{(M_{A},N_{A})}$ are drawn either from a GUE or GOE ensemble. While exact expressions for the first moments differ between GUE and GOE, they are asymptotically the same in the thermodynamic limit [39, 52, 44]. Fig. S1(a) confirms that there is little difference between the average entanglement values. This is consistent with the results found in Ref.[48]. Figure S1(b) also shows that the standard deviation of $\sigma_{M,N}$(GUE) is a factor $\sqrt{2}$ larger than $\sigma_{M,N}$(GOE) (shown by the blue dashed line). The additional prefactor is due to averaging over both the real and complex parts of the GUE ensemble.
@@
For the Hamiltonian in Eq.(Entanglement patterns of quantum chaotic Hamiltonians with a scalar U(1) charge) of the main text, when computing the first two moments of the EE distribution for mid-spectrum eigenstates, it is necessary to take a finite energy window $\Delta E$ in which to take samples of $S_{A}$, see Fig. S2(a). If the energy window is too small, then a statistically small number of states will be available for sampling, thus leading to large error bars. On the other hand, if the window is too large, then finite-energy eigenstates with low entanglement will skew the distribution and increase its variance, see Figs.S2(c). Because of typicality, we argue only a few eigenstates are necessary to quantify the mean and standard deviation. In Fig. S2 (a) we plot the eigenstate entanglement of the full spectrum for different system sizes, highlighting the set of eigenstates used in the main text to construct the microcanonical average [indicated with vertical lines in (b) and (c)].